# Motional entanglement in low-energy collisions near shape resonances

### v1

@@ Conclusions
We have analyzed motional entanglement generated upon the collision of two spinless particles in three-dimensional space. While standard scattering theory is based on plane waves, we have found a time-dependent treatment starting from separated wave packets with finite position and momentum uncertainties to be crucial to capture entanglement generation in general and the role of interferences in particular. Quantifying entanglement via the inverse single-particle purity after scattering, we find a linear relationship between entanglement and the collision cross section for sufficiently narrow initial momentum dispersion, in line with earlier findings for pure $s$-wave scattering and very weak collisions [23]. We predict scattering resonances to maximize entanglement generation, unless the momentum dispersion is much broader than the width of the resonance in which case the signature of the resonance is smeared out. While for momentum dispersions smaller than the resonance width, maximum entanglement is generated at the same collision energy at which the cross section peaks, this shifts to the energy which maximizes the Wigner time delay for momentum dispersions comparable with the resonance width. An important next step will be to extend our framework to particles with internal structure, in order to allow for the description of inelastic and reactive collisions. This will open the way to studying the role of entanglement in chemical reaction dynamics.
@@ Acknowledgments
Appendix A Methods and Formulation
The single-particle purity ($\mathcal{P}=1/\mathcal{K}$) is obtained as:
$\mathcal{P}(t)$ $=$ $\int d^{3}\vec{r}_{1}d^{3}\vec{r}_{2}d^{3}\vec{r}_{1}^{\,\prime}d^{3}\vec{r}_{2}^{\,\prime}\,\,\tilde{\Psi}^{*}(\vec{r}_{1},\vec{r}_{2}^{\,\prime},t)\tilde{\Psi}^{*}(\vec{r}_{1}^{\,\prime},\vec{r}_{2},t)\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)\tilde{\Psi}(\vec{r}_{1}^{\,\prime},\vec{r}_{2}^{\,\prime},t)$ (3)
$\tilde{\Psi}$ is the scattering wave-packet, which can be derived from the non-scattering wave-packet $\Psi$}. When ${\Psi}(\vec{r}_{1},\vec{r}_{2},t)$ is expanded in the plane-wave basis, its coefficients are the single-particle wavefunctions in momentum space $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ [Eq. (7)]. The plane-wave basis is free to rotate from the separate-particle coordinates, into CM and relative coordinates, using the fractional masses $\mu_{\alpha}$ of each particle ($\mu_{1}+\mu_{2}=1$):
$\Psi(\vec{r}_{1},\vec{r}_{2},t)$ $=$ $\int d^{3}\vec{P}\int d^{3}\vec{p}\,\,\psi_{01}(\mu_{1}\vec{P}+\vec{p},t)\psi_{02}(\mu_{2}\vec{P}-\vec{p},t)e^{i\vec{P}\cdot(\mu_{1}\vec{r}_{1}+\mu_{2}\vec{r}_{2})}e^{i\vec{p}\cdot(\vec{r}_{1}-\vec{r}_{2})}\qquad\text{no sctt}$ (4)
$\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$ $=$ $\int d^{3}\vec{P}\int d^{3}\vec{p}\,\,\psi_{01}(\mu_{1}\vec{P}+\vec{p},t)\psi_{02}(\mu_{2}\vec{P}-\vec{p},t)e^{i\vec{P}\cdot(\mu_{1}\vec{r}_{1}+\mu_{2}\vec{r}_{1})}f(\vec{p},\vec{r}_{1}-\vec{r}_{2})\quad\text{scatter}$ (5)
The scattering wave-packet $\tilde{\Psi}$ differs from $\Psi$} only in the relative motion $\vec{r}=\vec{r}_{1}-\vec{r}_{2}$. For plane-wave scattering, $e^{i\vec{p}\cdot\vec{r}}\xrightarrow{sctt}f(\vec{p},\vec{r})$, where $f(\vec{p},\vec{r})$ is the eigenfunction of the full scattering Hamiltonian $H_{0}+V(r)$. For $\tilde{\Psi}$ with finite momentum dispersion, the same replacement applied to each $\vec{p}$ component, while the coefficients of the plane-wave basis keep unchanged, assuming scattering is an adiabatic process. For any isotropic scattering potential $V(r)$ decaying no slower than $r^{-2}$, $f(\vec{p},\vec{r})$ can be expanded into partial wave components as:
@@ Appendix A Methods and Formulation
with $P_{l}$ the Legendre polynomials, and
$\psi_{l}(p,r)\xrightarrow{r\rightarrow\infty}\frac{1}{2pr}[e^{2i\delta_{l}(p)}e^{i(pr-\frac{l+1}{2}\pi)}+e^{-i(pr-\frac{l+1}{2}\pi)}]$
There exists a cut-off $L$ beyond which scattering phase-shifts $\delta_{l}(p)\rightarrow 0$ and radial wavefunction $\psi_{l}(p,r)\rightarrow j_{l}(pr)$ the first kind spherical Bessel function, when the effective centripetal barrier $\propto\frac{L(L+1)}{r^{2}}$] large enough to block $V(r)$. For example, for finite range potentials in Table 1, we set $L=6$ in the calculation.
The single-particle momentum distribution $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ ($\alpha=1,2$) are Gaussian functions ($\hbar=1$):
$\psi_{0\alpha}(\vec{p}_{\alpha},t)=\left[\frac{\det A}{(4\pi^{3})^{3}}\right]^{\frac{1}{4}}\exp\left[{-\frac{1}{2}(\vec{p}_{\alpha}-\vec{p}_{0\alpha})^{T}A(\vec{p}_{\alpha}-\vec{p}_{0\alpha})}\right]\exp\left[{-\frac{it}{2m}p_{\alpha}^{2}}\right]\exp\left[{i\vec{p}_{\alpha}\cdot\vec{q}_{0\alpha}}\right]$ (7)
We assume the particles to have identical mass $m$ and dispersion $A=\mathrm{diag}(a_{x},a_{y},a_{z})$, with coaxially arranged momenta and locations $\vec{p}_{01}=-\vec{p}_{02}={p}_{0}\hat{z}$, $\vec{q}_{01}=-\vec{q}_{02}=\frac{{q}_{0}}{2}\hat{z}$, so the scattering state $\tilde{\Psi}$ is separable in the CM and relative coordinates. The temporal evolution is parameterized by $\frac{t}{m}$, as shown in Fig. 1(b), where $\mathcal{K}$ shows an inverse linear correlation with the overlap $F=|\langle\Psi|\tilde{\Psi}\rangle|^{2}$, fitted respectively as 6.58F(10,10,20)+6.57-6.58\,F(10,10,20)+6.57, 12.07F(1,1,20)+8.59-12.07\,F(1,1,20)+8.59, and 6.72F(1,1,2)+4.77-6.72\,F(1,1,2)+4.77. Table. 2 gives the fit parameters for post-scattering entanglement $\mathcal{K}(t\to+\infty)$ to cross-sections $\sigma_{fit}$ in Figs. 2 and 3.
Table 2: Parameters of the fit Eq. (2) for the initial conditions in Figs. 2(b,c) and 3.
@@
For the numerical calculations, Eq. (3) is integrated using Gauss-Legendre quadrature, with normalization accuracy $\langle\tilde{\Psi}|\tilde{\Psi}\rangle-1<10^{-2}$. The stabilization of $\mathcal{K}$ is confirmed by integrating in each direction of $\vec{r}_{\alpha}$ using different orders of Gauss-Legendre quadrature. The saturated entanglement shown in Fig. 2 and 3 is calculated with time near $t/m\approx 20+q_{0}/(2p_{0})$.
Appendix B Momentum-Space Analysis
Since the wave packet is expanding and oscillating in real space but not in momentum space, on first glance it might seem easier to calculate the entanglement in momentum space [25, 23, 27], using scattering state $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$:
$\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)=\int d^{3}\vec{p}\,\,\psi_{01}\left[\mu_{1}(\vec{p}_{1}+\vec{p}_{2})+\vec{p},t\right]\psi_{02}\left[\mu_{2}(\vec{p}_{1}+\vec{p}_{2})-\vec{p},t\right]\,\mathcal{F}(\vec{p},\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2})\,$, (8)
where $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ ($\alpha=1,2$) are the Gaussian functions in Eq. (7), $\mathcal{F}(\vec{p},\vec{p}_{12})$ is the Fourier transform of $f(\vec{p},\vec{r})$ [Eq. (6)] and $\vec{p}_{12}=\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2}$ is the relative momentum after scattering. $\mathcal{F}(\vec{p},\vec{p}_{12})=\delta(\vec{p}-\vec{p}_{12})$ only when no scattering occurs $f(\vec{p},\vec{r})=e^{i\vec{p}\cdot\vec{r}}$]. However, otherwise,
@@ Appendix B Momentum-Space Analysis
The principal value terms $\textit{P}\,[$...] [43] have been incorrectly ignored by a semi-classical model that treats scattering as a “reflection” of relative momentum [26, 27], effectively assuming $|\vec{p}|=|\vec{p}_{12}|$. This assumption preserves the time-dependency as $exp\,{[it{(p_{1}^{2}+p_{2}^{2})}/{2m}]}$ in $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$ during the scattering, which ignores the interaction $V$ and results in time-independent $\mathcal{K}$ [25, 23, 26, 27]. Our calculation shows that neglecting $\textit{P}\,[$...] will compromise normalization $\langle\tilde{\Psi}|\tilde{\Psi}\rangle$, violate energy conservation $\langle\tilde{\Psi}|H|\tilde{\Psi}\rangle$, and and underestimate $\mathcal{K}$ up to a factor of 3 under certain initial conditions comparing to its correct asymptotic value $\mathcal{K}(t\to+\infty)$.
To properly integrate over $\vec{p}$ in Eq. (8) including $\textit{P}\,[$...], however, is challenging: The singularity near $p_{12}$ in $\textit{P}\,[$...] is usually handled by the Residue theorem, assuming the remaining integrand is analytic and vanishes at large distance. However, exponential factors such as Gaussian functions $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ diverge over half of the complex plane [e.g., $e^{-(iR)^{2}}\to+\infty$], blocking the construction of a closed contour $C$, making the Residue theorem no longer advantageous. Therefore, evaluating $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$ is not easy as initially appears, but rather tricky and difficult. To avoid these issues and fully capture coherence, we evaluate entanglement in real space through $\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$. Nevertheless, Eq. (9) remains instructive for elucidating why plane-wave initial states fail to generate entanglement, as we explain next.
Appendix C Why plane waves forbid entanglement generation
Assume that the initial state satisfies any two of the following three conditions: (1) $\Delta{P}=0$. (2) $\Delta{p}=0$. (3) The particles are separable prior to collision. Then the pre-scattering state takes the form:
@@ Appendix C Why plane waves forbid entanglement generation
where $N=(2\pi)^{3}\delta(\vec{0})$ is the normalization factor, $\vec{p}_{0\alpha}$ is the initial momentum for particle $\alpha$, and $\vec{p}_{0}=\mu_{2}\vec{p}_{01}-\mu_{1}\vec{p}_{02}$ and $\vec{P}_{0}=\vec{p}_{01}+\vec{p}_{02}$ are the relative and CM momentum. $E_{0}=\frac{p_{01}^{2}}{2m_{1}}+\frac{p_{02}^{2}}{2m_{2}}$ is the eigenenergy, the particle masses $m_{1,2}$ and $\mu_{i}=m_{i}/(m_{1}+m_{2})$ are arbitrary.
The scattering wave-packet $\tilde{\Psi}$ and the single-particle purity therefore become:
@@
$\mathcal{P}$ $=(2\pi)^{12}\int d^{3}\vec{p}_{1}d^{3}\vec{p}_{2}d^{3}\vec{p}_{1}^{\,\prime}d^{3}\vec{p}_{2}^{\,\prime}\,\,\tilde{\Psi}^{*}(\vec{p}_{1},\vec{p}_{2}^{\,\prime},t)\tilde{\Psi}^{*}(\vec{p}_{1}^{\,\prime},\vec{p}_{2},t)\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)\tilde{\Psi}(\vec{p}_{1}^{\,\prime},\vec{p}_{2}^{\,\prime},t)=\frac{(2\pi)^{9}}{N^{3}}\int d^{3}\vec{p}_{12}|\mathcal{F}(\vec{p}_{0},\vec{p}_{12})|^{4}\,.$ (12)
where $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})$ is given in Eq. (9), and can be decompose into non-scattered and a scattered components: $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})=\delta(\vec{p}_{0}-\vec{p}_{12})+\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})$, with the latter including finite terms of partial waves characterized by $e^{2i\delta_{l}(p_{0})}-1]$. For $\vec{p}_{0}=\vec{p}_{12}$, $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})=\delta(\vec{p}_{0}-\vec{p}_{12})\gg\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})$, otherwise $\mathcal{F}(\vec{p}_{0},\vec{p}_{12})=\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})$. Therefore, separating the integral $\int d^{3}\vec{p}_{12}$ into contributions for $\vec{p}_{12}=\vec{p}_{0}$ and $\vec{p}_{12}\neq\vec{p}_{0}$ yields
$\begin{split}\mathcal{P}&=1+\frac{(2\pi)^{9}}{N^{3}}\int_{\vec{p}_{12}\neq\vec{p}_{0}}d^{3}\vec{p}_{12}|\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p}_{12})|^{4}\geq 1\,.\end{split}$ (13)

### v2

@@ Conclusions
We have analyzed motional entanglement generated upon the collision of two spinless particles. Collisions in one- and three-dimensional space are strikingly different, and this reflects also in the generation of entanglement. While standard scattering theory is based on plane waves, we have found a time-dependent treatment starting from separated wave packets with finite position and momentum uncertainties to be crucial: It avoids violation of energy conservation and captures the generation of entanglement, including the role of interferences correctly. Quantifying entanglement via the inverse single-particle purity after scattering, we find a linear relationship between the asymptotic entanglement and the collision cross section for sufficiently narrow initial momentum dispersion. We predict that scattering resonances maximize entanglement generation, unless the momentum dispersion is much broader than the width of the resonance. As the momentum dispersion increases, becoming comparable to the resonance width, the collision energy that maximizes entanglement shifts from the peak of the cross section to the peak of the Wigner time delay. Future work should extend this framework to particles with internal structure, enabling investigations into entanglement swapping across diverse degrees of freedom and the generation of multi-partite entanglement in collisions. Furthermore, generalizing this approach to inelastic and reactive collisions will provide a quantitative assessment of the role of entanglement in chemical reaction dynamics, a widely conjectured connection [Devolder2020, Devolder2021, devolder2025] that has yet to be quantified. Finally, our work highlights the need for establishing concrete protocols for experimental measurement of the entanglement in collisions, as this will be essential for validating any theoretical prediction.
@@ Acknowledgments
End Matter
Calculation of the single-particle purity
The single-particle purity $\mathcal{P}=1/\mathcal{K}=Tr(\rho_{1}^{2})$ is obtained as
$\mathcal{P}(t)$ $=$ $\int d^{3}{r}_{1}d^{3}{r}_{2}d^{3}{r}_{1}^{\,\prime}d^{3}{r}_{2}^{\,\prime}\,\,\tilde{\Psi}^{*}(\vec{r}_{1},\vec{r}_{2}^{\,\prime},t)\tilde{\Psi}^{*}(\vec{r}_{1}^{\,\prime},\vec{r}_{2},t)\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)\tilde{\Psi}(\vec{r}_{1}^{\,\prime},\vec{r}_{2}^{\,\prime},t)\,$, (3)
where $\rho_{1}(\vec{r}_{1},\vec{r}_{1}^{\,\prime},t)=\int d^{3}r^{\,\prime}_{2}\,\tilde{\Psi}^{*}(\vec{r}_{1},\vec{r}_{2}^{\,\prime},t)\tilde{\Psi}(\vec{r}_{1}^{\,\prime},\vec{r}_{2}^{\,\prime},t)$ and $\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$ is the scattered wave packet. It differs from the unscattered wave packet $\Psi(\vec{r}_{1},\vec{r}_{2},t)$ only in the relative motion $\vec{r}=\vec{r}_{1}-\vec{r}_{2}$. Expanding $\Psi(\vec{r}_{1},\vec{r}_{2},t)$ into the plane-wave basis with expansion coefficients $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ and rotating from separate particle coordinates into CM and relative coordinates (using fractional masses $\mu_{\alpha}$ with $\mu_{1}+\mu_{2}=1$),
$\Psi(\vec{r}_{1},\vec{r}_{2},t)$ $=$ $\int d^{3}{P}\int d^{3}{p}\,\,\psi_{01}(\mu_{1}\vec{P}+\vec{p},t)\psi_{02}(\mu_{2}\vec{P}-\vec{p},t)e^{i\vec{P}\cdot(\mu_{1}\vec{r}_{1}+\mu_{2}\vec{r}_{2})}e^{i\vec{p}\cdot(\vec{r}_{1}-\vec{r}_{2})}$
$\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$ $=$ $\int d^{3}{P}\int d^{3}{p}\,\,\psi_{01}(\mu_{1}\vec{P}+\vec{p},t)\psi_{02}(\mu_{2}\vec{P}-\vec{p},t)e^{i\vec{P}\cdot(\mu_{1}\vec{r}_{1}+\mu_{2}\vec{r}_{2})}f(\vec{p},\vec{r}_{1}-\vec{r}_{2})$
where $f(\vec{p},\vec{r})$ are the eigenfunctions of the full scattering Hamiltonian $H_{0}+V(r)$. Assuming that scattering is an adiabatic process, the expansion coefficients remain unchanged upon scattering. For any isotropic potential $V(r)$ decaying no slower than $1/r^{2}$, $f(\vec{p},\vec{r})$ can be expanded into partial wave components,
@@
where $P_{l}$ is a Legendre polynomial and
$\psi_{l}(p,r)\xrightarrow{r\rightarrow\infty}\frac{1}{2pr}[e^{2i\delta_{l}(p)}e^{i(pr-\frac{l+1}{2}\pi)}+e^{-i(pr-\frac{l+1}{2}\pi)}]\,.$
For large $l$, $\delta_{l}(p)\rightarrow 0$ due to the high centrifugal barrier; for the finite range potentials in Table 1, the cutoff $L=6$ is sufficient in the calculations.
We take the single-particle momentum distributions $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ ($\alpha=1,2$) to be Gaussians (with $\hbar=1$),
$\psi_{0\alpha}(\vec{p}_{\alpha},t)=\left[\frac{\det A}{(4\pi^{3})^{3}}\right]^{\frac{1}{4}}\exp\left[{-\frac{1}{2}(\vec{p}_{\alpha}-\vec{p}_{0\alpha})^{T}A(\vec{p}_{\alpha}-\vec{p}_{0\alpha})}\right]\exp\left[{-\frac{it}{2m}p_{\alpha}^{2}}\right]\exp\left[{i\vec{p}_{\alpha}\cdot\vec{q}_{0\alpha}}\right]\,$, (5)
assuming identical mass $m$ and dispersion $A=\mathrm{diag}(a_{x},a_{y},a_{z})$, such that the scattered state $\tilde{\Psi}$ is separable in the CM and relative coordinates. We integrate Eq. (3) using Gauss-Legendre quadratures along each direction of $\vec{r}_{\alpha}$. $\mathcal{K}$ is stabilized over the choice of mesh with at least two significant digits, with normalization accuracy $\langle\tilde{\Psi}|\tilde{\Psi}\rangle-1<10^{-2}$. The post-scattering entanglement $\mathcal{K}(t/m)$ shows an inverse linear correlation with the overlap $F=|\langle\Psi|\tilde{\Psi}\rangle|^{2}$ as shown in Fig. 1(b), fitted as 6.58F+6.57-6.58\,F+6.57, 12.07F+8.59-12.07\,F+8.59, and 6.72F+4.77-6.72\,F+4.77 for $\vec{a}=(10,10,20)$, $(1,1,20)$ and $(1,1,2)$, respectively. The saturated entanglement $\mathcal{K}(t\to+\infty)$ is calculated at $t/m\approx 20+q_{0}/(2\,p_{0})$, and the fit parameters for the cross-sections $\sigma_{fit}$ in Figs. 2 and 3 are given in Table 2.
Table 2: Parameters of the fit $\sigma_{fit}(\mathcal{K})=\gamma(\mathcal{K}-\beta)$ for the initial conditions in Figs. 2(b,c) and 3.
@@
Momentum-Space Analysis
Since the wave packet is expanding and oscillating in real space but not in momentum space, on first glance it might seem easier to calculate the entanglement in momentum space [Tal2005PRL, WangPRA2006, Harshman2008, Weder2011, WEDER201394],
$\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)=\int d^{3}{p}\,\,\psi_{01}\left[\mu_{1}(\vec{p}_{1}+\vec{p}_{2})+\vec{p},t\right]\psi_{02}\left[\mu_{2}(\vec{p}_{1}+\vec{p}_{2})-\vec{p},t\right]\,\mathcal{F}(\vec{p},\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2})\,$, (6)
where $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ ($\alpha=1,2$) are the Gaussians from Eq. (5), $\mathcal{F}(\vec{p},\vec{p}_{12})$ is the Fourier transform of $f(\vec{p},\vec{r})$, cf. Eq. (4), and $\vec{p}_{12}=\mu_{2}\vec{p}_{1}-\mu_{1}\vec{p}_{2}$ is the relative momentum after scattering. Importantly, $\mathcal{F}(\vec{p},\vec{p}_{12})=\delta(\vec{p}-\vec{p}_{12})$ only when no scattering occurs, otherwise
@@
The principal value term $\textit{P}\,[$...] appears in both 1D and 3D scattering [Dirac1927], but is ignored when treating scattering as a “reflection” of relative momentum [Schmuser2006, Harshman2008], which effectively assumes $|\vec{p}|=|\vec{p}_{12}|$. This assumption results in a time-dependency of $\exp\,{[it{(p_{1}^{2}+p_{2}^{2})}/{2m}]}$ in $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$, which yields a time-independent $\mathcal{K}_{s}$. However, according to our calculations, $\mathcal{K}_{s}$ can be very different from the correct asymptotic value $\mathcal{K}(t\to+\infty)$  [^3] Neglecting the time-dependence may also explain claims of $\mathcal{K}_{max}$ larger than 2 in 1D collisions, as entanglement overall decays with time in 1D, converging to the bound expected for two orthogonal states.. Moreover, neglecting $\textit{P}\,[$...] results in a violation of energy conservation $\langle\tilde{\Psi}|H|\tilde{\Psi}\rangle$. It also eliminates the relative phase-shifts between the incoming and outgoing partial wave components, i.e., a key feature of quantum scattering.
It is challenging to properly integrate over $\vec{p}$ in Eq. (6): The singularity near $p_{12}$ in $\textit{P}\,[$...] is usually handled by the residue theorem, assuming the remaining integrand is analytic and vanishes at large distance. However, exponential factors as those in the Gaussians $\psi_{0\alpha}(\vec{p}_{\alpha},t)$ diverge over half of the complex plane, impeding the common construction of a closed contour $C$. Moreover, the analyticity of the phase-shifts $\delta_{l}(p)$ over the complex plane can be problematic. All of this makes evaluating $\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)$ rather tricky, and entanglement calculations based on momentum space need to be carefully scrutinized. To avoid these issues and fully capture coherence, we have evaluated entanglement in real space through $\tilde{\Psi}(\vec{r}_{1},\vec{r}_{2},t)$, as described above. Nevertheless, Eq. (7) remains instructive for elucidating why plane-wave initial states fail to generate entanglement, as we explain next.
Why plane waves preclude the generation of entanglement
Assume that the initial state satisfies any two of the following three conditions: (1) $\Delta{P}=0$. (2) $\Delta{p}=0$. (3) The particles are separable prior to collision. Then the pre-scattering state takes the form
@@
where $N$ is a normalization factor, $\vec{p}_{0\alpha}$ is the initial momentum for particle $\alpha$, and $\vec{p}_{0}=\mu_{2}\vec{p}_{01}-\mu_{1}\vec{p}_{02}$ and $\vec{P}_{0}=\vec{p}_{01}+\vec{p}_{02}$ are the relative and CM initial momenta, $E_{0}={p_{01}^{2}}/{(2m_{1})}+{p_{02}^{2}}/{(2m_{2})}$ is the scattering energy (here, we make no assumptions on the particle masses $m_{1,2}$). The scattering wave-packet $\tilde{\Psi}$ and the single-particle purity then become
@@
$\mathcal{P}$ $=(2\pi)^{12}\int d^{3}{p}_{1}d^{3}{p}_{2}d^{3}{p}_{1}^{\,\prime}d^{3}{p}_{2}^{\,\prime}\,\,\tilde{\Psi}^{*}(\vec{p}_{1},\vec{p}_{2}^{\,\prime},t)\tilde{\Psi}^{*}(\vec{p}_{1}^{\,\prime},\vec{p}_{2},t)\tilde{\Psi}(\vec{p}_{1},\vec{p}_{2},t)\tilde{\Psi}(\vec{p}_{1}^{\,\prime},\vec{p}_{2}^{\,\prime},t)=\frac{(2\pi)^{9}}{N^{3}}\int d^{3}\vec{p}\,|\mathcal{F}(\vec{p}_{0},\vec{p})|^{4}\,$, (10)
where $\mathcal{F}(\vec{p}_{0},\vec{p})$, cf. Eq. (7), can be decomposed into non-scattered and a scattered components, $\mathcal{F}(\vec{p}_{0},\vec{p})=\delta(\vec{p}_{0}-\vec{p})+\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p})$ with
[…]
For $\vec{p}_{0}=\vec{p}$, the first term dominates, $\mathcal{F}(\vec{p}_{0},\vec{p})=\delta(\vec{p}_{0}-\vec{p})\gg\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p})$, as $\delta_{l\to\infty}(p_{0})\to 0$. Otherwise $\mathcal{F}(\vec{p}_{0},\vec{p})=\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p})$. Therefore, separating the integral $\int d^{3}\vec{p}$ into contributions for $\vec{p}=\vec{p}_{0}$ and $\vec{p}\neq\vec{p}_{0}$ yields
$\begin{split}\mathcal{P}&=1+\frac{(2\pi)^{9}}{N^{3}}\int_{\vec{p}\neq\vec{p}_{0}}d^{3}\vec{p}\,|\tilde{\mathcal{F}}(\vec{p}_{0},\vec{p})|^{4}\geq 1\,.\end{split}$ (12)